OpenStudy (noahbred):

i will medal What effort force would you apply to a screwdriver to pry open a lid with a resistance of 1,000 N? The screwdriver has an MA of 50. (Remember to rearrange the formula.) 0.05 N 20 N 200 N 50,000 N

OpenStudy (anonymous):

\[mechanical~advantage=\frac{ output~force }{ input~force }\]\[MA=\frac{ F_{out} }{ F_{inn} }\]

OpenStudy (anonymous):

\[Input~force=?\]\[Output~force=1000N\]\[MA=50\]

OpenStudy (anonymous):

Substitute the value into the formula

OpenStudy (anonymous):

\[50=\frac{ 1000 }{ F_{input} }\]

OpenStudy (anonymous):

\[F_{input}=\frac{ 1000 }{ 50 }\]

OpenStudy (anonymous):

\[F_{input}=?\]solve it @noahbred

OpenStudy (noahbred):

1000n

OpenStudy (pawanyadav):

F (input)=1000/50=100/5=?

OpenStudy (pawanyadav):

What we get when divided 100 by 5?

OpenStudy (pawanyadav):

It is 20 ...right